OpenStudy (anonymous):

What's the positive square root of 0.81?

OpenStudy (yanasidlinskiy):

.9

OpenStudy (anonymous):

Thanks honey!

OpenStudy (yanasidlinskiy):

No problem.

OpenStudy (phi):

one way to do this if you don't want to use a calculator is notice that 0.81 is the same as \[ \frac{81}{100} \] to find the square root of a fraction you can find the square root of its top and its bottom, then simplify \[ \frac{\sqrt{81}}{\sqrt{100}}= \frac{9}{10} \]